OpenStudy (anonymous):

The minute-hand of a clock moves 1/10 degree every second. If you look at the clock when the minute hand is past the 12,you can use the equation 1=1/10x+10 to find how many degrees past the 12 minute hand after x seconds. Graph the equation and tell whether it is linear.

OpenStudy (anonymous):

I think this is really just asking you to graph the equation y = (1/10)x + 10 If so, it is like graphing any other equation you've seen that looks like y = mx + b

OpenStudy (anonymous):

I get that part, But what should go in the x part?

OpenStudy (anonymous):

I think there is a typo in the equation you listed... the part on the left side should be the y, I think. x is the number of seconds that have gone by. So after 60 seconds, you would have y = (1/10)(60) + 10 = 60/10 + 10 = 6 + 10 = 16

OpenStudy (anonymous):

That point would be plotted as (60, 16). You can plot other points the same way by picking out different numbers of seconds for the x value.

OpenStudy (anonymous):

My bad. Yeah, It says y=1/4x-1

OpenStudy (anonymous):

ok, well, different equation then but same idea. Have you done stuff like y = mx + b, slope, y-intercept, and that sort of material? If so, you can look at that equation and tell it's a line with slope 1/4 and y intercept at y = -1.

OpenStudy (anonymous):

So I can do like 60,65,and 70? &Yup.

OpenStudy (anonymous):

Otherwise, just pick x values... like 60, 65, 70, etc , and solve for the y values using your equation, then plot those points as ordered pairs (x,y)

OpenStudy (anonymous):

Okay. Thanks. :)

OpenStudy (anonymous):

It should be a line, though... just as a heads-up, but you need to graph it to show that you know that. But if you start getting a weird zig-zaggy line, check your points, 'cause you probably made an error on one of them.

OpenStudy (anonymous):

Okay.